a person completes half of his journey with a speed of v1 and other half with speed of v2. the average speed of person is ?
 

Grade:11

1 Answers

Upendra Tagare
6 years ago
AVG speed = total distance / total timet1=0.5s/v1 and t2=0.5s/V2AVG speed = s/(t1+t2)AVG speed = 2v1v2/(v1+V2)